OpenStudy (anonymous):

How do you find the maximum height of an object on it's parabolic path with three points?

OpenStudy (amistre64):

simply :)

OpenStudy (anonymous):

How is that?

OpenStudy (amistre64):

all points are equi distant from the foci; we can use that, to determine a proper equation for the ellipse

OpenStudy (amistre64):

of course, we have to consider that this forms a circle as well; which is a kind of special ellipse

OpenStudy (amistre64):

parabolic ..... did you change that from ellipse?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

if so, thats even simpler

OpenStudy (amistre64):

there are a variety of methods that would work; do you have a preference?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

hmmm, do you have 3 points to work with so that its not so absract at least?

OpenStudy (anonymous):

yes...

OpenStudy (amistre64):

really making me pull teeath on this one aintcha .....

OpenStudy (anonymous):

(18,8) (45,0) (0,1.39)

OpenStudy (amistre64):

well, the y = 0, and the x=0 points are useful for one of the simpler methods

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

im going with an old standby; developing an equation from successives zeros take the x values given (0,45,18) and create some zeros with them: (0-x)(45-x)(18-x) set up an equation such that we zero out all but one variable y = a + bx + cx(45-x) when x=0, this reduces to: y = a, or simply 1.39 = a let x = 45, and y = 0 0 = 1.39 + b(45) + c45(45-45) -1.39 = b(45) ; b = -1.39/45 this gives us so far y = 1.39 + -1.39x/45 + cx(45-x) so let x = 18 to solve for c

OpenStudy (amistre64):

8 = 1.39 -1.39(18)/45 + c18(45-18) 8 - 1.39 +1.39(18)/45 = c18(45-18) (8 - 1.39 +1.39(18)/45)/(18(45-18)) = c therefore .... y = 1.39 -1.39x/45 + (8 - 1.39 +1.39(18)/45)x(45-x)/(18(45-18)) is the equation :) the wolf can simplify that as needed http://www.wolframalpha.com/input/?i=y+%3D++1.39+-1.39x%2F45+%2B+%288+-+1.39+%2B1.39%2818%29%2F45%29x%2845-x%29%2F%2818%2845-18%29%29

OpenStudy (anonymous):

thanks so much :)

OpenStudy (amistre64):

there is a matrix method as well, but in the end .... we get the same results :) good luck